The profit of an organization is calculated by the function P(x)=x^2-4000x + 7800000 where x is the number of units sold. If the net profit 3800000, find the number of items sold.
Answer by lwsshak3(11628)

The profit of an organization is calculated by the function P(x)=x^2-4000x + 7800000 where x is the number of units sold. If the net profit 3800000, find the number of items sold.
***
P(x)=x^2-4000x + 7800000
3800000=x^2-4000x + 7800000
x^2-4000x+4000000=0
(x-2000)^2=0
x=2000
P(2000)=3800000
number of items sold=2000